The inverse demand curve for a monopolist’s product is P = -2Q + 20. At which price is the
elasticity of demand equal to −1?
1
Expert's answer
2017-03-03T10:29:05-0500
Firstly, we need find equal Q, that`s why P = -2Q + 20, and we get Q= -0.5P+10, and E p = b*P/Q, where b=-0.5, E p =-1. That`s why -1= -0.5*P/Q, and P/Q=2, where P/(-0.5P+10)=2, 2P=20, P=10

ANSWER: P=10